OpenStudy (anonymous):

-5x^2+11x+6 Is this prime? And if it is why it is.

OpenStudy (anonymous):

It doesn't seem to factor with rational coefficients; you can check by solving for the roots of the function. If the roots are rational, you can usually factor.\[-5x^2+11x+6=0 \implies x=\frac{11 \pm \sqrt{241}}{10}\]Irrational roots means not factorable.

OpenStudy (anonymous):

So that means it is prime, correct?

OpenStudy (anonymous):

Yup. Do math every day.

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

No sweat.